OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = 3x + 4, what is f(g(x)) ?

OpenStudy (anonymous):

12x + 16 - 1 = 12x -15

OpenStudy (angela210793):

f[g(x)]=fog=4(3x+4)-1=12x+15

OpenStudy (anonymous):

4(3x+4)-1 12x+16-1

OpenStudy (lalaly):

4(3x+4)-1 = 12x+16-1 = 12x+15

OpenStudy (anonymous):

12x + 15 typo

OpenStudy (anonymous):

x = 3x + 4 4(3x+4) - 1 12x + 16 - 1 12x + 15

OpenStudy (anonymous):

thank you all so much!!!